# A Hybrid Quantum-Classical Particle-in-Cell Method for Plasma Simulations

### v1

@@ 1 Introduction
    We use a physics-informed neural network approach with a Poisson PDE loss to improve the accuracy of predictions.
@@ 3.2 Experimental Setup
The two-stream instability problem that we use to train the model and validate the results has the following parameters. The simulation is initialized with two counter-streaming beams of electrons at velocities $v_{0}$ (red points in phase-space) and $-v_{0}$ (blue points in phase-space). We consider the maximum value of $v_{0}=0.1$ for experimentation that results in the instability in the plasma system. The thermal velocity $v_{th}$ is set to 0. The particles move within a domain of length $L=1$. The PIC simulation set-up has 64 grid points, and each point represents 200 particles. The total simulation time is 50, which is split into 1000 steps (step size = 0.05), and in each step, the PIC cycle is performed to update the particle movement. All the physical quantities discussed in the paper are dimensionless, with the electron charge to mass ratio equal to 1.
Figure 3: The left figure shows the electric fields $\mathrm{E}=-\mathrm{\nabla\Phi}$ prediction for some of the time-steps of the PIC simulation for $v_{0}=0.07$. The predictions by the CQC, and CCC PIC codes are compared with the traditional baseline PIC code. The right figure shows the mean relative absolute error (MRAE) in predicting the electric fields for 1000 time steps.
@@
The classical layer that we use is always a linear layer (implemented with torch.nn.Linear) with input and output features set to 64, as we need to encode data at 64 grid points. The classical layer is always followed by a ReLU activation function layer (except for the last layer, where we use Tanh activation function). The quantum layer consists of a data encoding part and a trainable part as shown in Fig. 1. We use amplitude encoding scheme to embed the incoming data from the classical layer into the quantum circuit. In this encoding scheme, $2^{n}$ features are encoded in the probability amplitude of $n$ qubit state (As $n$ qubits are described by a vector of dimension $2^{n}$ in the Hilbert space). Therefore, in our model, the real number outputs from the classical layers (of dimensionality 64) are encoded using 6 qubits. The quantum circuit is initialized in the state,
@@
where the probability amplitudes are normalized, $\sum\limits_{i}|w_{i}|^{2}=1$.
The trainable part of the quantum circuit consists of the rotation (Rot) gates and controlled-NOT (CNOT) gates. A rotation gate takes 3 arguments, which are trainable. The role of CNOT gate is to entangle different qubits. We have used $\mathrm{NL}=6$ such trainable layers within a quantum layer of the hybrid model. Further, we have experimented with different read-outs of the quantum circuit, which are passed to the subsequent neural network layers. They can be quantum states (the real part of the quantum state amplitudes), probabilities of different computational basis states, and the expectation value of the Pauli Z operator measurement. We found out that the probabilities resulted in a better PIC performance and is a practically implementable choice.
@@ 3.2.1 Training
We simulate the 1D 2-stream instability problem with the traditional linear equation solvers to compute electric potentials for 1000 time steps. We use 500 instances from this data set as the training data. The neural network receives the charge density at 64 grid-cell points as the input, and there are 64 features in the output, which should approximate the electric potential at those grid points when the parameters of the model are optimized using a loss function.
@@
where $\lambda$ is the relative weight between data loss and physics loss. We have used the Adam optimizer with a learning rate of 0.001 for all our experiments. The neural networks are optimized in 2000 epochs. We have also preprocessed the inputs ($\rho$) and output dataset ($\Phi$), by simply dividing by the maximum of absolute values (maximum taken over $x$) at each time step,
@@ 4.1 Data-driven supervised learning approach
Here we present the results of using the supervised data-driven approach, i.e, the loss function (MAE) is evaluated at the 64 grid points using the given true electric potential points, as described in Eq. (4). The training data is generated from a baseline PIC code with initial stream velocities $v_{0}=[0.03,0.05,0.1]$. The tests are made on the PIC simulation with a new unseen stream velocity $v_{0}=0.07$, which should reveal the generalizing capacity of the neural networks under consideration.
@@
In Figure 3, we present the electric field prediction results for different PIC iterations. The left figure shows the electric fields $E(x)$ predictions for some of the time steps when compared with the ones obtained using traditional methods ($\Phi_{exact}$). We observe that the hybrid CQC model can fit the data well, while the classical data deviates from the correct solutions in some of the instances. The plots also demonstrate the highly non-linear nature of the electric fields in question and the diversity in their shapes, which the hybrid neural network is able to capture efficiently. We compare the overall predictions for the 1000 time steps of the simulations in the boxplot shown on the right in Figure 3. The plot shows the distribution of the mean relative absolute errors, MRAE (the relative metric is chosen because the electric field amplitudes lie within a large range of values), where the error values by a CQC model has a lower median value ($\approx 0.168$) and the smaller outliers than the classical CCC model (median $\approx 0.265$).
@@
Further, in Figure 4, we show the phase space of the particles at the end of the simulation for $v_{0}=0.07$. The instability results in a vortex-like structure in the phase space of the particles, which is characteristic of the 2-stream instability problem. One can see that both CQC and CCC models result in a phase space similar to the baseline PIC simulation, however, a careful observation of the velocity distribution $f(v)$ as shown in the histogram reveals that CQC leads to a phase space that is closer to the traditional PIC simulation.
@@
We gain further insights into the problem by comparing the total energy of the plasma throughout simulations, as shown in Figure 5. The total energy of the particles during the simulation should be conserved, and the numerical value should be $\approx 0.0024$. While in the initial phase, both models conserve energy; at later times, the total energy is not conserved. We see that the CQC-PIC code shows an energy fluctuation that is slightly smaller when compared to the energy fluctuations of the CCC-PIC code. In addition, we also track the electric field amplitudes during the evolution. The electric field amplitudes are similar to the traditional PIC simulation, with the exponential growth of the electric field in the initial phase and the saturation of amplitudes when the instability is fully developed [20].
@@ 4.2 Physics-informed learning approach
In this section, we present the results obtained using physics-informed neural networks integrated with PIC code for the 2-stream instability problem. In this case, the neural networks are trained with the loss function $\mathcal{L}_{\mathrm{PINN}}$ in Eq. (5). We observe improved accuracy in predicting electric fields, particularly in the case of classical neural networks, which previously yielded relatively inaccurate results under the purely data-driven approach, as discussed in the previous section.
We use $N_{d}=64$ data points and the Poisson PDE loss evaluated at $N_{g}=64$ points in Eq. (5). The gradients are computed using torch.gradient() module. We use a relative weight $\lambda=0.3$ (for CCC) and $\lambda=0.5$ (for CQC) between data loss and physics loss, and the models are trained for 3000 epochs. In Fig. 6 we present the results of errors in predicting true electric fields for the case of $v_{0}=0.07$, where the errors are computed against the true solutions obtained from the traditional methods. We compare the performance of both the CQC and CCC neural networks when trained with $\mathcal{L}_{\mathrm{PINN}}$ and $\mathcal{L}_{\mathrm{data}}$. The physics-informed approach shows an advantage for the classical CCC model in predicting the electric field; however, when the trained model is integrated into the PIC iterations, there is no particular advantage in phase space. For the hybrid model, the improvement is marginal in electric field predictions, and the resultant PIC evolution showed slightly lower fluctuations in total energy when compared to the data-driven approach. Additionally, the PINN method should yield reliable results even with a reduced number of data points $N_{d}$, which is particularly advantageous in practical scenarios where only sparse data is available for training the neural network.
@@
Furthermore, we have used the finite difference method (which is implemented using torch.gradient()) to compute the gradients within the physics loss as we are dealing with the discrete set of $\Phi$ values of the neural network output. Using more efficient methods, such as automatic differentiation techniques to define $\mathcal{L}_{\mathrm{PINN}}$, can improve the physics-informed strategy and make it more effective.
@@
We test the physics-informed HNN approach for another special case of 2-streams of electrons, where the two beams are initialized with a normal distribution of thermal velocities $v_{th}$ with mean equal to 0. In this case, theoretically, there should not be any growth in the induced electric field, and the plasma should remain stable. The total energy is also conserved in such a system. The CCC and CQC models are trained with the data of $v_{th}=[0.01,0.03,0.07,0.1]$. We make the test hybrid PIC simulation for the case of $v_{th}=0.05$. In Figure 7, we show the results of total energy and the final velocity distribution. Both CQC and CCC lead to small energy fluctuations, with CQC showing lower fluctuation than CCC. Both the classical and hybrid models result in a similar velocity distribution at the end of the simulation and match well with that of the traditional PIC simulation.
@@ 4.3 Parallelization, Performance and Computational Cost
We show the training results of CQC model using multiple MPI processes in Figure 8. The data parallelization helps in reducing the time required to complete a single epoch. However, the averaging of loss leads to slow convergence of the loss function over the number of epochs. In Figure 8, first, we show the loss reduction for 1000 epochs using the MPI processes from 1 to 16. At the end of 1000 epochs, different processes have reached different stages of training. In the second plot, we show the tradeoff in the number of processes and the loss minimization after 1000 steps. The plot indicates that using 6 MPI processes works best in terms of both (among the number of processes we have experimented with). A CQC model takes $\approx 700s$ to complete 1000 epochs (the training data size is 500) of training, while CCC model takes $\approx 37s$. A PIC simulation with CCC and CQC as Poisson solvers takes $55s$ and $28.5s$ respectively for 1000 PIC iterations. The CQC model has $8,428$ trainable parameters, and the CCC model has $12,480$ parameters. It is to be noted that a CQC model has fewer trainable parameters, which leads to reliable results, which could be a potential advantage when scaled to large models applied to more complicated problems.
Figure 8: Results of MPI-based data-parallelization strategy. The first plot shows the training of CQC network using different number of processes. The second plot shows the trade-off curve in speedup and loss function minimization.
5 Related Work
Several works have been proposed to solve the governing equations of plasma governing equations. Refs. [21, 7, 22] presents an overview of the different approaches for the solution of plasmas equation in the context of fusion energy and plasma physics, including Hamiltonian simulation approaches and quantum lattice algorithms (QLA) [23].
The work presented in this paper is a hybrid approach based on the solution of the Poisson equation using quantum neural networks (QNNs) and a scientific quantum machine learning approach. The properties of QNNs have been studied in a number of different works, focusing on expressivity, trainability, and power of generalization [11, 24, 25]. QNNs have been used for solving a number of different applications, ranging from computational fluid dynamics [26, 14, 17] to condensed matter physics [27].
QNNs have also been used for solving PDEs, for example, through physics- [14, 15, 28]. Previous works, such as Refs. [29, 15, 16], focus on using QNNs for solving the Poisson equation. In these works, quantum neural networks were trained to solve the given instance of the Poisson equation. Integrating such unsupervised methods within the PIC loop can be computationally expensive, as they require a long time to train. In this paper, we use a supervised learning approach, where we train HNNs to solve the Poisson equation that can also generalize to unseen instances. A similar strategy was also adopted in Ref. [30], where the authors introduced supervised deep neural networks and deep convolutional networks to solve the Poisson equation within PIC codes. Moreover, in several of the recent works [31, 32, 26], classical-quantum hybrid neural networks have been utilized to perform machine learning tasks, which have shown advantages over their classical counterparts when it comes to accuracy, mostly likely due to the high expressive power of the quantum circuits [11]. Motivated by these recent works, we integrate the classical-quantum hybrid neural networks within traditional PIC codes as the Poisson equation solvers.

### v2

@@ 1 Introduction
    We use a physics-informed neural network approach with a Poisson PDE loss to improve the accuracy of predictions, which works particularly well in regimes with sparse data availability.
@@ 2 Background
3 Related Work
Several quantum computing methods have been proposed to solve the governing equations of plasma. Refs. [19, 7, 20] presents an overview of the different strategies plasma simulation in the context of fusion energy and plasma physics, including Hamiltonian simulation approaches and quantum lattice algorithms (QLA) [21].
Quantum algorithms, in particular linear solvers, have also been proposed to solve discretized PDEs, such as the Poisson equation [22, 23]. Most of these approaches fall within the category of fault-tolerant quantum algorithms. Our work is based on the variational approaches using quantum neural networks (QNNs). The properties of QNNs have been studied in a number of different works, focusing on expressivity, trainability, and power of generalization [11, 24, 25].They have been employed to solve problems within varying fields, ranging from computational fluid dynamics [26, 14, 17] to condensed matter physics [27].
QNNs have also been used for solving PDEs, for example, through physics-informed neural networks [14, 15, 28]. Previous works, such as Refs. [29, 15, 16], focus on using QNNs for solving the Poisson equation. In these works, quantum neural networks were trained to solve the given instance of the Poisson equation. Integrating such unsupervised methods within the PIC loop can be computationally expensive, as they require a long time to train. In this paper, we use a supervised learning approach, where we train HNNs to solve the Poisson equation that can also generalize to unseen instances. A similar strategy was also adopted in Ref. [30], where the authors introduced supervised deep neural networks and deep convolutional networks to solve the Poisson equation within PIC codes. Moreover, in several of the recent works [31, 32, 26], classical-quantum hybrid neural networks have been utilized to perform machine learning tasks, which have shown advantages over their classical counterparts when it comes to accuracy, most likely due to the high expressive power of the quantum circuits [11]. Motivated by these recent works, we integrate the HNNs within traditional PIC codes as the Poisson equation solvers.
Figure 3: The left figure shows the electric fields $\mathrm{E}=-\mathrm{\nabla\Phi}$ prediction for some of the time-steps of the PIC simulation for $v_{0}=0.07$. The predictions by the CQC, and CCC PIC codes are compared with the traditional baseline PIC code. The midddle figure shows the mean relative absolute error (MRAE) in predicting the electric fields for 1000 time steps. The right figure shows the median of MRAEs of predicting the electric fields by the HNN models with varying number of quantum trainable parameters (Q params in the figure). The performance is compared with the CCC models with a similar number of trainable parameters.
@@ 4.2 Experimental Setup
The two-stream instability problem that we use to train the model and validate the results has the following parameters. The simulation is initialized with two counter-streaming beams of electrons at velocities $v_{0}$ (red points in phase-space) and $v_{0}$ (blue points in phase-space). We consider the maximum value of $v_{0}=0.1$ for experimentation. This initial velocity is sufficient to result in the instability within the plasma system. The thermal velocity $v_{th}$ is set to 0. The particles move within a domain of length $L=1$. The PIC simulation set-up has 64 grid points, and each point represents 200 particles. The total simulation time is 50, which is split into 1000 steps (step size = 0.05), and in each step, the PIC cycle is performed to update the particle movement. All the physical quantities discussed in the paper are dimensionless, with the electron charge to mass ratio equal to 1.
@@
The classical layer that we use is always a linear layer (implemented with torch.nn.Linear) with input and output features set to 64, as we need to encode data at 64 grid points. The classical layer is always followed by an ReLU activation function layer (except for the last layer, where we use Tanh activation function).
The quantum layer is implemented using a quantum circuit. This layer essentially has three major components as shown in Fig. 1: a data encoding part (sometimes also referred to as a quantum feature map), a variational ansatz block and measurements of the quantum circuit outputs. Here we briefly explain the implementation of each of these crucial parts.
Quantum Embedding
We use amplitude encoding scheme to embed the incoming data from the classical layer into the quantum circuit. In this encoding scheme, $2^{n}$ features are encoded in the probability amplitude of $n$ qubit state (As $n$ qubits are described by a vector of dimension $2^{n}$ in the Hilbert space). Therefore, in our model, the real number outputs from the first classical layer (of dimensionality 64) are encoded using 6 qubits. The quantum circuit is initialized in the state,
@@
where the probability amplitudes are normalized, $\sum\limits_{i}|w_{i}|^{2}=1$. We implement this encoding using Pennylane’s qml.AmplitudeEmebdding tool. Here, the quantum circuit is prepared in the desired initial state using a sequence of uniformly controlled rotations as proposed by Möttönen et.al. [34]. This scheme requires $2^{n+2}-4n-4$ CNOT gates and $2^{n+2}-5$ single qubit rotation gates for the state preparation.
Variational ansatz
This is the trainable part of the quantum circuit that generally consists of the rotation gates and controlled-Unitary gates. In this work, we use Pennylane’s Strongly entangling layers (qml.StronglyEntanglingLayer) template as the variational ansatz. It consists of general rotation (Rot) gate that takes 3 arguments, which are trainable. This is followed by several Controlled NOT (CNOT) gates. The role of CNOT gate is to entangle different qubits. This particular variational ansatz has been used extensively in the literature [24, 31, 35, 36] and has shown to yield accurate results. We have used the number of trainable layers $\mathrm{NL}=6$ to present our main results. Further, we have tested the usage of other two ansatz layers that are defined in Pennylane. One is qml.BasicEntanglerLayer which consists of single parameter rotation gates (RX gates) and CNOT gates [24]. The other ansatz is the simplified two-design layers proposed in [37] (qml.SimplifiedTwoDesign) which consists of RY gates (a single parameter rotation gate) and entangling CZ gates. We discuss the performance of each variational ansatz in detail in Sec. 5.1.
Measurements
The qubits undergo a unitary transformation determined by the variational ansatz described previously and the final state of the qubits is subsequently measured. We have experimented with different read-outs of the quantum circuit, which are passed to the following classical neural network layers. The measurement outcome of the quantum circuit can be the quantum state itself. Although the statevector simulators can return quantum state description, on a real quantum computer, advanced quantum state tomography techniques have to be used [38]. Alternatively, one can compute the Born rule probability distribution of measurements in the computational basis and in practice this requires reinitializing and measuring the same quantum circuit in multiple shots. One can also compute the expectation value of an observable from such probability distributions, which turns out to be the most commonly used quantum output in variational quantum algorithms. After comparing the different strategies discussed above, we found that using computational basis measurement probability distributions, which is also easier to implement in practice, resulted in better accuracy.
Figure 4: Comparison of the performance of different numbers of layers for the three variational ansatz in HNNs: Basic Entangling layer, Simplified Two-design and Strongly entangling layers. The number of trainable Quantum parameters in each model considered is mentioned on top of the bars.
@@ 4.2.1 Training
We simulate the 1D 2-stream instability problem with the traditional linear equation solvers to compute electric potentials for 1000 time steps. We use 500 instances from this dataset as the training data. The HNN receives the charge density at 64 grid-cell points as the input, and there are 64 features in the output. The output should approximate the electric potential at the 64 grid points when the parameters of the model are optimized using an appropriate loss function.
@@
where $\lambda$ is the relative weight between data loss and physics loss. $N_{d}$ is the number of data points used to compute the data-loss in the PINN approach. The gradients are computed using torch.gradient() module.
We have used the Adam optimizer with a learning rate of 0.001 for all our experiments. We have also preprocessed the inputs ($\rho$) and output dataset ($\Phi$), by simply dividing by the maximum of absolute values (maximum taken over $x$) at each time step,
@@ 5.1 Data-driven supervised learning approach
Here we present the results of using the supervised data-driven approach, i.e, we use the loss function in Eq. (4) evaluated at the $N_{g}=64$ grid points. The training data is generated from a baseline PIC code with initial stream velocities $v_{0}=[0.03,0.05,0.1]$. Here, the CQC model used has 8,428 trainable parameters, and the CCC model has 12,480 parameters. Both models are trained for 2000 epochs. The tests are made on the PIC simulation with a new unseen stream velocity $v_{0}=0.07$, which should reveal the generalizing capacity of the neural networks under consideration.
@@
In Figure 3, we present the electric field prediction results for different PIC iterations. The left figure shows the electric fields $E(x)$ predictions for some of the time steps when compared with the ones obtained using traditional methods ($\Phi_{exact}$). We observe that the hybrid CQC model can fit the data well, while the CCC prediction deviates from the correct solutions in some of the instances. The plots also demonstrate the highly non-linear nature of the electric fields in question and the diversity in their shapes, which the HNN is able to capture efficiently. We compare the overall predictions for the 1000 time steps of the simulations in the boxplot shown on the right in Fig 3. The plot shows the distribution of the mean relative absolute errors, MRAE (the relative metric is chosen because the electric field amplitudes lie within a large range of values), where the error values by a CQC model has a lower median value ($\approx 0.168$) and the smaller outliers than the classical CCC model (median $\approx 0.265$). Furthermore, in Fig. 3 we evaluate the HNN model using different numbers of variational layers (Strongly entangling layers), each corresponding to a different count of trainable parameters, and report the median of MRAEs for electric field predictions. Our results show a decreasing trend in prediction error as the number of quantum parameters increases. We also compare the HNN’s performance to that of the CCC model with a similar number of trainable parameters and observe that, within this range, the HNN consistently outperforms the classical model.
Additionally, we evaluate the performance of various variational ansatz described in Sec.4.2. Specifically, we consider three circuit structures: Strongly Entangling Layers, Basic Entangling Layers, and a simplified Two-Design ansatz. For each ansatz, we vary the number of layers, $NL=[2,4,6,8,10]$, and present the corresponding results in Fig.4. Among the three, the Strongly Entangling Layers consistently achieve the best performance. This advantage is likely due to the presence of three-parameter single-qubit rotations in the ansatz, which introduce more quantum parameters for a given number of layers. The number of quantum parameters is indicated above each bar in the plot. The number of classical trainable parameters is fixed at 8192 across all configurations. However, it is not evident to what extent the quantum properties such as entanglement is contributing to the high expressive nature of the quantum circuits and requires further investigation.
@@
We finally integrate the trained HNN with the PIC code and present the HNN-PIC simulation results in Fig. 5, where we show the phase space of the particles at the end of the simulation for $v_{0}=0.07$. The instability results in a vortex-like structure in the phase space of the particles, which is characteristic of the 2-stream instability problem. One can see that both CQC and CCC models result in a phase space similar to the baseline PIC simulation. However, the velocity distribution $f(v)$ as shown in the histogram reveals that CQC leads to a phase space that is closer to the traditional PIC simulation. We computed the energy distance (a statistical test to compare the similarity between two probability distributions as described in [39], implemented using the python module scipy.stats.energy_distance) between the predicted velocity distribution and the baseline PIC simulation results. The lower the energy distance, the more similar the distributions are to each other. The energy distance between CQC PIC and baseline PIC was found to be $0.012$, and that between CCC and baseline PIC was found to be $0.016$. This test confirms that the CQC PIC simulated phase space is closer to the one simulated by the traditional PIC method in terms of the energy distance measure.
@@
We gain further insights into the problem by comparing the total energy of the plasma throughout simulations, as shown in Fig 6. The total energy of the particles during the simulation should be conserved, and the numerical value should be $\approx 0.0024$. While in the initial phase, both models conserve energy; at later times, the total energy is not conserved. We see that the CQC-PIC code shows an energy fluctuation that is slightly smaller when compared to the energy fluctuations of the CCC-PIC code. In addition, we also track the electric field amplitudes during the evolution. The electric field amplitudes are similar to the traditional PIC simulation, with the exponential growth of the electric field in the initial phase and the saturation of amplitudes when the instability is fully developed [40].
@@ 5.2 Physics-informed learning approach
In this section, we present the results obtained using physics-informed neural networks integrated with PIC code for the 2-stream instability problem. In this case, the neural networks are trained with the loss function $\mathcal{L}_{\mathrm{PINN}}$ in Eq. (5). All the models considered for the experimentation in this section are trained for 3000 epochs.
Physics-informed approaches are known to produce reliable results even in the presence of sparse data. Therefore, here we have considered two scenarios when the training data (electric potential $\Phi$) is dense, in other words, has a higher spatial resolution, and another case when the training data has solutions at sparse points. This test can potentially mimic the situation of low grid resolution data from the experiments or simulations to train the neural networks with. We train the CCC and CQC models with both data-driven and physics-informed approaches. We compare the errors (MRAE) in predicting the electric field ($\mathrm{E}=-\mathrm{\nabla\Phi}$) for the case of $v_{0}=0.07$ when the models are trained with the data for the cases $v_{0}=[0.03,0.05,0.1]$.
As the first test, we use $N_{d}=64$ data points (same as the data-driven strategy) to compute the data loss $\mathcal{L}_{\mathrm{data}}$ and the Poisson PDE loss $\mathcal{L}_{\mathrm{PINN}}$ evaluated at $N_{g}=64$ points in Eq. (5). We use a relative weight $\lambda=0.3$ (for CCC) and $\lambda=0.5$ (for CQC) between data loss and physics loss. In Fig. 7, the first four boxplots show the prediction results with the dense data scenario. We observe a marginal improvement in the CCC model when using the PINN approach. In the case of CQC, PINN results in negligible improvements. This is likely due to the fact that the performance of the neural networks is saturated due to the good quality of data which also obey the correct PDEs. Therefore, the physics-informed approach is not further enhancing the prediction accuracy.
As the second test, we consider sparse points of the electric potential $N_{d}=20$ in the training dataset. The physics loss is computed for $N_{g}=64$ collocation points. The prediction results for the case of sparse data are shown in the last four boxplots in Fig. 7. Purely data-driven approaches using both HNNs and classical networks perform poorly in generalising to predict electric fields for the test case. However, we observe a drastic improvement when using PINN approach for both CCC and CQC models in comparison to the respective purely data-driven approaches. Furthermore, sparse-data PINNs perform with errors on par with dense-data models. We have also confirmed that the improvement in the performance by the PINN method is statistically significant by performing Wilcoxon signed-rank test [41] between different pairs of error distributions depicted in Fig. 7, and we have obtained p-values <0.01<0.01 in each case.
@@
Moreover, we have used the finite difference method (which is implemented using torch.gradient()) to compute the gradients within the physics loss as we are dealing with the discrete set of $\Phi$ values of the neural network output. Using more efficient methods, such as automatic differentiation techniques to define $\mathcal{L}_{\mathrm{PINN}}$, can improve the physics-informed strategy and make it more effective.
@@
We test the physics-informed HNN approach for another special case of 2-streams of electrons, where the two beams are initialized with a normal distribution of thermal velocities $v_{th}$ with mean equal to 0. In this case, theoretically, there should not be any growth in the induced electric field, and the plasma should remain stable. The total energy is also conserved in such a system. The CCC and CQC models are trained with the data of $v_{th}=[0.01,0.03,0.07,0.1]$. We make a test hybrid PIC simulation for the case of $v_{th}=0.05$. In Fig. 8, we show the results of total energy and the final velocity distribution. Both CQC and CCC lead to small energy fluctuations, with CQC showing lower fluctuation than CCC. We further study the velocity distribution at the end of the simulation in each case. The energy distance computed between CQC-PINN and traditional $f(v)$ is 0.00061, and that between CCC-PINN and traditional $f(v)$ is 0.00059. The energy distance values and the histogram shown in Fig. 8 reveal that both the classical and hybrid models result in a similar velocity distribution at the end of the simulation and match well with that of the traditional PIC simulation.
@@ 5.3 Parallelization, Performance and Computational Cost
We show the training results of CQC model using multiple MPI processes in Fig. 9. The data parallelization helps in reducing the time required to complete a single epoch. However, the averaging of loss leads to slow convergence of the loss function over the number of epochs. In Fig. 9, first, we show the loss reduction for 1000 epochs using the MPI processes from 1 to 16. At the end of 1000 epochs, different processes have reached different stages of training. In the second plot, we show the tradeoff in the number of processes and the loss minimization after 1000 steps. The plot indicates that using 6 MPI processes works best in terms of both (among the number of processes we have experimented with). A CQC model takes $\approx 700s$ to complete 1000 epochs (the training data size is 500) of training, while CCC model takes $\approx 37s$. A PIC simulation with CCC and CQC as Poisson solvers takes $55s$ and $28.5s$ respectively for 1000 PIC iterations. The CQC model has $8,428$ trainable parameters, and the CCC model has $12,480$ parameters. It is to be noted that a CQC model has fewer trainable parameters, which leads to reliable results, which could be a potential advantage when scaled to large models applied to more complicated problems.
Figure 9: Results of MPI-based data-parallelization strategy. The first plot shows the training of CQC network using different numbers of processes. The second plot shows the trade-off curve between speedup and loss function minimization.